X Company makes two products, A and B, and uses an activity-based costing overhead allocation system, with three cost pools and three cost drivers. Budgeted costs and driver information for 2020 were as follows:

Cost Pool 1 Cost Pool 2 Cost Pool 3
Budgeted costs $110,000     $51,000     $33,000    
Cost Drivers
  Product A 81,400     56,700     71,000    
  Product B 72,700     49,700     45,300    


What will the allocation to Product A be? [round overhead rate(s) to two decimal places].
Answer #1

Overhead rate = cost / cost drivers

1 = 110000/(81400+72700) = $0.71

2 = 51000 /(56700+49700) = $0.48

3 = 33000 /(71000+45300) = $0.28

Allocation to producr A = 81400*0.71 + 56700*0.48 + 71000*0.28

= $104890